Grade 12th passAlgebra

Find the square of the following numbers.
86........

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

(86)2
= (90-4)2
= (90)2 + (4)2 - 2×90×4 [Since, (a+b)2 = a2+b2 +2ab]
= 8100+16–720
= 8116–720
= 7396

Thanks